# August 2025 APCTP Pre2025-017

### v1

@@ 1 Introduction
In supersymmetry, the inflaton scalar should be assigned to an irreducible massive $\mathcal{N}=1$ supermultiplet that can be either a chiral one or a vector one. If the inflaton multiplet is chiral (i.e. described by a chiral superfield), its physical scalar field component must be complex, i.e. real inflaton must be accompanied by the pseudo-scalar superpartner called sinflaton. The scalar potential is given by the so-called $F$-term, where $F$ is the auxiliary scalar of the chiral superfield and an order parameter of supersymmetry breaking. We employ the $F$-term scalar potential in this paper, without using gauge fields. [^1] A massive vector supermultiplet has a real massive physical scalar that is suitable as the inflaton for embedding some single-field inflation models whose scalar potential is given by a real function squared, known as the $D$-term, where $D$ is the auxiliary field of the vector supermultiplet and the order parameter of supersymmetry breaking as well [25, 26, 27]. Multi-field inflation with multiple or non-Abelian vector supermultiplets is beyond the scope of this work.
@@ 2.2 Shift symmetry and stabiliser superfield
However, this also implies that $K_{i}$ depends only on the imaginary part that vanishes or is suppressed in many simple models of inflation. Then the large-field behaviour of the potential looks like $V\sim K^{\bar{j}i}W_{i}\bar{W}_{\bar{j}}-3|W|^{2}$, where for generic (e.g., polynomial) $W(\Phi)$ the second term will dominate.222 This can be evaded by exponential superpotentials. To obtain a sufficiently flat potential, however, a delicate balance between multiple exponential terms is required [49, 50, 51, 52]. See also Refs. [53, 54, 55] for more recent developments. As a result, the potential becomes negative in the large-field region, namely, $V\sim-3|W|^{2}<0$.
@@ 2.3 Shift-symmetric higher-order terms
Let us consider a generic $W$ with the following $K$:
@@
where $c$ and $\xi$ are real parameters. The quadratic term gives the approximately canonical kinetic term. The quartic term ensures the imaginary part of $\Phi$ approximately vanishes as long as $\Phi$ breaks supersymmetry and $\xi>+\mathcal{O}(1)$ because it leads to supersymmetry-breaking mass squared term $12\xi(H^{2}+m_{3/2}^{2})$. Then, the first term leads to a non-vanishing $K_{\Phi}$, which boosts the positive contribution to the scalar potential. The absence of the cubic term can be understood as follows. Suppose there is initially a cubic term as well as other shift-symmetric terms. Provided that the quartic term is sufficiently large, the imaginary part is stabilised at a finite field value. Expanding $K$ around this point, the cubic term approximately vanishes.
@@ 3.5 More general Kähler potentials
with the real parameters $\alpha_{i}$ and $\xi_{i}$.555 The parameters $\alpha_{i}$ in front of the logarithm are inspired by the $\alpha$-attractor models [64, 65, 66, 67], but there is an important difference. In the $\alpha$-attractor-type or typical no-scale-type models, the inflaton is usually identified with a field combination appearing in the argument of the logarithm. Here, we identify the orthogonal direction in the field space with the inflaton, which is protected by the manifest shift symmetry. Our $\alpha_{i}$’s are similar to the parameters $c_{i}$ introduced above, as can be seen after Taylor expansion of the logarithm. In fact, the expressions below (after stabilisation of the imaginary parts) depend only on the combination $\alpha_{i}+c_{i}$, where $c_{i}$’s are the parameters in the superpotential $W=e^{ic_{i}\Phi^{i}}\widetilde{W}$.
@@ 4 Stabilisation mechanism for multi-superfield inflation
We always assume a shift-symmetric $K=K(\Phi^{i}-\bar{\Phi}^{\bar{i}})$. As in the single-superfield case, we neglect the cubic terms in KK.666 Effects of the residual cubic term as well as the higher-dimensional terms were studied in Ref. [38] for the single-superfield case. We first briefly explain the underlying reason for this. The general shift-symmetric $K$, after canonically normalising the quadratic term for simplicity of the analyses, can be written as
@@
where the subscript $\mathcal{G}$ denotes the (time-dependent) goldstino component, e.g., $\tau_{i\mathcal{G}\mathcal{G}}\equiv\tau_{ijk}f^{j}f^{k}$ and $\xi_{ij\mathcal{G}\mathcal{G}}\equiv\xi_{ijkl}f^{k}f^{l}$, with $f^{i}\equiv\langle G^{i}/\sqrt{G_{j}G^{j}}\rangle$ being the goldstino direction, $H$ is the Hubble parameter, $m_{3/2}=e^{G/2}$ is the gravitino mass, and we used the Friedmann equation. As can be seen from the above equation, if $\hat{\tau}_{ijk}$’s are too large, the minimum of the imaginary parts is significantly deviated from the origin. Therefore, expanding $K$ around the minimum, the cubic terms should be negligible. [^7] Precisely speaking, the linear term $\hat{\kappa}_{i}$ also shifts the minimum, which is more carefully studied in Subsec. 4.1 Thus, we assume that there is a point in the field space where the cubic terms are negligible. Around the point, or in terms of the shifted field $I^{i}\simeq\hat{I}^{i}-(\hat{\xi}^{-1})^{ij\mathcal{G}\mathcal{G}}\hat{\tau}_{j\mathcal{G}\mathcal{G}}/2\sqrt{2}$ where $\hat{\xi}^{-1}$ is the inverse matrix of $\hat{\xi}$ focusing on the $(i,j)$ indices, the cubic terms are approximately absent. Explicitly performing this shift will be difficult unless one knows all the higher-dimensional terms. The residual constant ($K_{0}$) and linear ($ic_{i}I^{i}$) terms can be transferred to the superpotential by the super-Weyl-Kähler transformation. After canonical normalisation of the quadratic term, $K$ becomes of the form
@@
This result can be understood as the goldstino-direction component ($\mathcal{G}=f_{i}\Phi^{i}$) of $\xi_{ijkl}$, i.e., $\mathcal{K}(f)\simeq 2\xi_{\mathcal{G}\mathcal{G}\mathcal{G}\mathcal{G}}=2\xi_{ijkl}f^{i}f^{j}f^{k}f^{l}$ (see Appendix A for details), whose sign depends upon a convention used in the literature.888 It is also instructive to note in the single-superfield case that the sign flips between the sectional curvature and the Ricci scalar in the field space as $\mathcal{R}=K^{\Phi\bar{\Phi}}K^{\Phi\bar{\Phi}}\mathcal{R}_{\Phi\bar{\Phi}\Phi\bar{\Phi}}=-\mathcal{K}(\Phi)$. For instance, the sectional curvature for the $\alpha$-attractor [64, 65, 66, 67] in our notation is $2/(3\alpha)$, where $\alpha=1$ corresponds to the no-scale models. The Kähler manifolds of the $\alpha$-attractor models have hyperbolic geometry, and the hyperbolic geometry has the negative sectional curvature in our notation. On the other hand, we should require $\mathcal{K}(f)$ to be positive for the stabilisation of the sinflaton.
@@ 5 Conclusion
Let us emphasise again the significance of our results and locate them in the context of inflation in supergravity. First, our proposal is not just specific models but a mechanism that can accommodate many inflation models. Second, there are only several known mechanisms that embed generic inflation models into supergravity such as $D$-term inflation models [69, 70, 71, 72, 73, 74, 75, 25, 26, 27] and models with stabiliser fields (original models mainly characterised by superpotential [28, 29, 30, 31] and those by Kähler potential [76, 77] (see also their cousins [78, 79, 80, 81, 82, 83, 84, 85]). Thus, our proposal significantly broadens the model-building possibilities for inflation in supergravity.
@@ Acknowledgements
This can be viewed as part of a unitary rotation in field space,999 We require the leading (unit matrix) part of the Kähler metric to be invariant under such transformation, i.e., $\delta_{i\bar{j}}=U^{\dagger}{}^{-1}_{\bar{j}}{}^{\bar{l}}\delta_{k\bar{l}}U^{-1}{}^{k}{}_{i}$. It follows $U^{-1}{}^{i}{}_{j}=\delta^{i\bar{k}}U^{\dagger}_{\bar{k}}{}^{\bar{l}}\delta_{j\bar{l}}$. which is also holomorphic,  [^10] The components of $U$ are to be regarded as complex numbers rather than fields because of the slow-roll assumption. Under the slow-roll assumption, we can also neglect the derivatives acting on the transformation matrix arising from the kinetic term. i.e. $\widehat{\Phi}^{i}=U^{i}{}_{j}\Phi^{j}$ (this also implies $\bar{\widehat{\Phi}}^{\bar{i}}=\bar{\Phi}^{\bar{j}}U^{\dagger}{}_{\bar{j}}{}^{\bar{i}}$), with
@@
    As regards the Planck bound [3] on isocurvature perturbations, by using Ref. [87], the standard constraint [88, 89, 90] reads

### v2

# December 2025 APCTP Pre2025-017

@@ 1 Introduction
In supersymmetry, the inflaton scalar should be assigned to an irreducible massive $\mathcal{N}=1$ supermultiplet that can be either a chiral one or a vector one. If the inflaton multiplet is chiral (i.e. described by a chiral superfield), its physical scalar field component must be complex, i.e. real inflaton must be accompanied by the pseudo-scalar superpartner called sinflaton (or vice versa). The scalar potential is given by the so-called $F$-term, where $F$ is the auxiliary scalar of the chiral superfield and an order parameter of supersymmetry breaking. We employ the $F$-term scalar potential in this paper, without using gauge fields. [^1] A massive vector supermultiplet has a real massive physical scalar that is suitable as the inflaton for embedding some single-field inflation models whose scalar potential is given by a real function squared, known as the $D$-term, where $D$ is the auxiliary field of the vector supermultiplet and the order parameter of supersymmetry breaking as well [25, 26, 27]. Multi-field inflation with multiple or non-Abelian vector supermultiplets is beyond the scope of this work.
@@ 2.2 Shift symmetry and stabiliser superfield
However, this also implies that $K_{i}$ depends only on the imaginary part that vanishes or is suppressed in many simple models of inflation. Then the large-field behaviour of the potential looks like $V\sim K^{\bar{j}i}W_{i}\bar{W}_{\bar{j}}-3|W|^{2}$, where for generic (e.g., polynomial) $W(\Phi)$ the second term will dominate. [^2] This can be evaded by exponential superpotentials. To obtain a sufficiently flat potential, however, a delicate balance between multiple exponential terms is required [49, 50, 51, 52]. See also Refs. [53, 54, 55] for more recent developments. As a result, the potential becomes negative in the large-field region, namely, $V\sim-3|W|^{2}<0$.
@@ 2.3 Shift-symmetric higher-order terms
Let us consider a generic $W$ with the following shift-symmetric $K$:
@@
where $c$ and $\xi$ are real parameters. The quadratic term gives the approximately canonical kinetic term. The quartic term ensures the imaginary part of $\Phi$ approximately vanishes as long as $\Phi$ breaks supersymmetry and $\xi>+\mathcal{O}(1)$ because it leads to supersymmetry-breaking mass squared term $12\xi(H^{2}+m_{3/2}^{2})$, where $H$ is the Hubble parameter and $m_{3/2}=e^{K/2}|W|$ is the gravitino mass. Then, the first term leads to a non-vanishing $K_{\Phi}$, which boosts the positive contribution to the scalar potential. The absence of the cubic term can be understood as follows. Suppose there is initially a cubic term as well as other shift-symmetric terms. Provided that the quartic term is sufficiently large, the imaginary part is stabilised at a finite field value. Expanding $K$ around this point, the cubic term approximately vanishes. Because of the shift symmetry in $K$, the flat direction $\text{Re}\,\Phi$ is the inflaton candidate, whose potential is specified by the superpotential $W$, which explicitly breaks the shift symmetry.
@@ 3.5 More general Kähler potentials
with the real parameters $\alpha_{i}$ and $\xi_{i}$. [^5] The parameters $\alpha_{i}$ in front of the logarithm are inspired by the $\alpha$-attractor models [64, 65, 66, 67], but there is an important difference. In the $\alpha$-attractor-type or typical no-scale-type models, the inflaton is usually identified with a field combination appearing in the argument of the logarithm. Here, we identify the orthogonal direction in the field space with the inflaton, which is protected by the manifest shift symmetry. Our $\alpha_{i}$’s are similar to the parameters $c_{i}$ introduced above, as can be seen after Taylor expansion of the logarithm. In fact, the expressions below (after stabilisation of the imaginary parts) depend only on the combination $\alpha_{i}+c_{i}$, where $c_{i}$’s are the parameters in the superpotential $W=e^{ic_{i}\Phi^{i}}\widetilde{W}$.
@@ 4 Stabilisation mechanism for multi-superfield inflation
We always assume a shift-symmetric $K=K(\Phi^{i}-\bar{\Phi}^{\bar{i}})$. As in the single-superfield case, we neglect the cubic terms in KK. [^6] Effects of the residual cubic term as well as the higher-dimensional terms were studied in Ref. [38] for the single-superfield case. We first briefly explain the underlying reason for this. The general shift-symmetric $K$, after canonically normalising the quadratic term for simplicity of the analyses, can be written as
@@
where the subscript $\mathcal{G}$ denotes the (time-dependent) goldstino component, e.g., $\tau_{i\mathcal{G}\mathcal{G}}\equiv\tau_{ijk}f^{j}f^{k}$ and $\xi_{ij\mathcal{G}\mathcal{G}}\equiv\xi_{ijkl}f^{k}f^{l}$, with $f^{i}\equiv\langle G^{i}/\sqrt{G_{j}G^{j}}\rangle$ being the goldstino direction, and we used the Friedmann equation. As can be seen from the above equation, if $\hat{\tau}_{ijk}$’s are too large, the minimum of the imaginary parts is significantly deviated from the origin. Therefore, expanding $K$ around the minimum, the cubic terms should be negligible. [^7] Precisely speaking, the linear term $\hat{\kappa}_{i}$ also shifts the minimum, which is more carefully studied in Subsec. 4.1 Thus, we assume that there is a point in the field space where the cubic terms are negligible. Around the point, or in terms of the shifted field $I^{i}\simeq\hat{I}^{i}-(\hat{\xi}^{-1})^{ij\mathcal{G}\mathcal{G}}\hat{\tau}_{j\mathcal{G}\mathcal{G}}/2\sqrt{2}$ where $\hat{\xi}^{-1}$ is the inverse matrix of $\hat{\xi}$ focusing on the $(i,j)$ indices, the cubic terms are approximately absent. Explicitly performing this shift will be difficult unless one knows all the higher-dimensional terms. The residual constant ($K_{0}$) and linear ($ic_{i}I^{i}$) terms can be transferred to the superpotential by the super-Weyl-Kähler transformation. After canonical normalisation of the quadratic term, $K$ becomes of the form
@@
This result can be understood as the goldstino-direction component ($\mathcal{G}=f_{i}\Phi^{i}$) of $\xi_{ijkl}$, i.e., $\mathcal{K}(f)\simeq 2\xi_{\mathcal{G}\mathcal{G}\mathcal{G}\mathcal{G}}=2\xi_{ijkl}f^{i}f^{j}f^{k}f^{l}$ (see Appendix A for details), whose sign depends upon a convention used in the literature. [^8] It is also instructive to note in the single-superfield case that the sign flips between the sectional curvature and the Ricci scalar in the field space as $\mathcal{R}=K^{\Phi\bar{\Phi}}K^{\Phi\bar{\Phi}}\mathcal{R}_{\Phi\bar{\Phi}\Phi\bar{\Phi}}=-\mathcal{K}(\Phi)$. For instance, the sectional curvature for the $\alpha$-attractor [64, 65, 66, 67] in our notation is $2/(3\alpha)$, where $\alpha=1$ corresponds to the no-scale models. The Kähler manifolds of the $\alpha$-attractor models have hyperbolic geometry, and the hyperbolic geometry has the negative sectional curvature in our notation. On the other hand, we should require $\mathcal{K}(f)$ to be positive for the stabilisation of the sinflaton.
@@ 5 Conclusion
Let us emphasise again the significance of our results and locate them in the context of inflation in supergravity. First, our proposal is not just specific models but a mechanism that can accommodate many inflation models. Second, there are only several known mechanisms that embed generic inflation models into supergravity such as $D$-term inflation models [69, 70, 71, 72, 73, 74, 75, 25, 26, 27] and models with stabiliser fields (original models mainly characterised by superpotential [28, 29, 30, 31] and those by Kähler potential [76, 77] (see also their cousins [78, 79, 80, 81, 82, 83, 84, 85])). Thus, our proposal significantly broadens the model-building possibilities for inflation in supergravity.
@@ Acknowledgements
This can be viewed as part of a unitary rotation in field space, [^9] We require the leading (unit matrix) part of the Kähler metric to be invariant under such transformation, i.e., $\delta_{i\bar{j}}=U^{\dagger}{}^{-1}_{\bar{j}}{}^{\bar{l}}\delta_{k\bar{l}}U^{-1}{}^{k}{}_{i}$. It follows $U^{-1}{}^{i}{}_{j}=\delta^{i\bar{k}}U^{\dagger}_{\bar{k}}{}^{\bar{l}}\delta_{j\bar{l}}$. which is also holomorphic,  [^10] The components of $U$ are to be regarded as complex numbers rather than fields because of the slow-roll assumption. Under the slow-roll assumption, we can also neglect the derivatives acting on the transformation matrix arising from the kinetic term. i.e. $\widehat{\Phi}^{i}=U^{i}{}_{j}\Phi^{j}$ (this also implies $\bar{\widehat{\Phi}}^{\bar{i}}=\bar{\Phi}^{\bar{j}}U^{\dagger}{}_{\bar{j}}{}^{\bar{i}}$), with
@@
    As regards the Planck bound [3] on the isocurvature perturbations of cold dark matter, by using Ref. [87], the standard constraint [88, 89, 90] reads